OpenStudy (anonymous):

solve for x (WILL FAN AND MEDAL!!!) -4x(+16)=116

OpenStudy (photon336):

\[-4x+16=116\] this is what you have right? @carsonce

OpenStudy (anonymous):

yep

OpenStudy (photon336):

Subtract 16 from both sides -4x+16 = 116-16 -16 This is what we get. -4x = 100 -4 -4 now divide both sides by -4 x = 100/-4 = -25 x = -25

OpenStudy (anonymous):

I just need to know how to identify the solution

OpenStudy (anonymous):

Oh wow thank you

OpenStudy (photon336):

@carsonce please let me know if there's something you don't get

OpenStudy (photon336):

we can check this too watch

OpenStudy (photon336):

-4(-25) = 100 + 16 = 116 116 = 116 this is how we check

OpenStudy (anonymous):

Ohhh that makes sense

OpenStudy (photon336):

if you need any more help let me know

OpenStudy (anonymous):

How would you do this? The perimeter of a rectangular photo is 72 cm. The length is twice the width. What are the length and width of the photo?

OpenStudy (anonymous):

@Photon336

OpenStudy (photon336):

let's translate this the length is TWICE the width So the length is 2x and the width is x Does this make sense?